Calculate the force of gravitation between the Earth and the Sun is given that the mass of the earth is equal to 6 x 10 raise to 24 kg and of the sun is equal to 2 x 10 raise to 30 kg the average distance between the two is 1.5 x 10 raise to 20


1
Expert's answer
2021-10-18T11:06:52-0400

We can find the force of gravitation as follows:


"F=\dfrac{GM_SM_E}{r^2},""F=\dfrac{6.67\cdot10^{-11}\ \dfrac{N\cdot m^2}{kg^2}\cdot2\cdot10^{30}\ kg\cdot6\cdot10^{24}\ kg}{(1.5\cdot10^{11}\ m)^2}=3.5\cdot10^{22}\ N."